OpenStudy (anonymous):

for the pair of functions,f(x)=x^2+2 and g(x)=x+5, find the following (a).(f*g)(x) (b).(f*g)(4) (c).(g*f)(x) (d).(g*f)(4)

OpenStudy (anonymous):

(f*g)x=f(x+5)= (x+5)^2 + 5 = x^2+10x+25+5 =x^2+10x+30 (f*g)(4) = 4^2+10*4+30 = 86

OpenStudy (anonymous):

(g*f)(x) =x^2 +2+5 = x^2+7 (g*f)(4)= 4^2+7 = 23